# MAVERIX: Multimodal Audio-Visual Evaluation Reasoning IndeX

### v1

@@ 3.2 Baselines
All models receive inputs in the unified format [video frames, subtitles, question], with frames uniformly sampled at their maximum supported context window. For audio-incapable models, subtitles replace raw audio tracks, while Gemini 1.5 Pro processes synchronized audio-video pairs. We employ a standardized prompt template across models, ensuring fairness by eliminating instructional biases. This setup isolates modality interdependence as the critical challenge: models must synthesize visual cues like the startled expression of a character with an auditory context like an off-screen crash to match human-like reasoning.
3.3 Results
Models Params Easy (%) Medium (%) Hard (%) Overall (%)
w/o sub w/ sub w/o sub w/ sub w/o sub w/ sub w/o sub w/ sub
Random Guess - 12.4 12.7 12.3 12.5
Human (Video) - 60.2 48.8 59.4 55.3
Human (Audio) - 34.9 27.4 40.8 31.1
Human (Video+Audio) - 83.6 76.2 84.4 80.3
Open-source MLLMs (Truncated Videos)
InternVL2 [14] 2B 31.1 35.2 (+4.1) 27.7 35.1 (+7.4) 17.8 24.3 (+6.5) 27.5 33.4 (+5.9)
Qwen2.5-VL [6] 7B 30.8 34.9 (+4.1) 25.9 31.9 (+6.0) 19.3 20.7 (+1.4) 26.8 31.3 (+4.5)
LLaVA-OneVision [27] 7B 26.6 28.9 (+2.3) 26.5 28.2 (+1.7) 29.9 29.9 (+0.0) 27.1 28.8 (+1.7)
DeepSeek-VL2-Tiny [51] 3.37B 16.8 15.1 (-1.7) 14.5 20.6 (+6.1) 20.5 17.3 (-3.2) 16.4 17.9 (+1.5)
DeepSeek-VL2-Small [51] 16.1B 39.9 46.4 (+6.5) 38.8 46.4 (+7.6) 25.9 29.6 (+3.7) 37.2 43.8 (+6.6)
Closed-source MLLMs (Truncated Videos)
GPT-4o [39] - 58.9 69.8 (+10.9) 60.7 67.0 (+6.3) 41.5 49.6 (+8.1) 56.9 65.4 (+8.5)
o1 [38] - 74.0 79.3 (+5.3) 70.7 76.0 (+5.3) 55.6 67.4 (+11.8) 69.6 75.9 (+6.3)
Claude Sonnet 3.5 [5] - 27.4 68.9 (+41.5) 39.6 64.6 (+25.0) 27.4 44.4 (+17.0) 40.3 63.1 (+22.8)
Grok2-Vision [52] - 58.3 65.1 (+6.8) 55.1 61.2 (+6.1) 45.9 47.4 (+1.5) 54.9 60.6 (+5.7)
Nova-Lite [23] - 51.8 59.8 (+8.0) 49.3 55.9 (+6.6) 37.8 43.0 (+5.2) 48.5 55.4 (+6.9)
Nova-Pro [23] - 53.0 51.5 (-1.5) 51.7 47.5 (-4.2) 35.6 34.1 (-1.5) 49.6 46.9 (-2.7)
Gemini 1.5 Pro [49] - 64.5 69.5 (+5.0) 65.4 69.4 (+4.0) 55.6 49.6 (-6.0) 63.5 66.3 (+2.8)
Open-source MLLMs (Full-Length Videos)
InternVL2 [14] 2B 23.0 26.3 (+3.3) 18.9 24.4 (+5.5) 14.2 11.0 (-3.2) 19.7 23.0 (+3.3)
Qwen2.5-VL [6] 7B 47.4 55.9 (+8.5) 45.3 51.7 (+6.4) 37.9 47.2 (+9.3) 44.8 52.6 (+7.8)
LLaVA-OneVision [27] 7B 34.5 40.1 (+5.6) 34.3 39.5 (+5.2) 30.7 35.4 (+4.7) 33.8 39.1 (+5.3)
DeepSeek-VL2-Tiny [51] 3.37B 17.1 26.0 (+8.9) 15.4 21.5 (+6.1) 21.3 19.7 (-1.6) 17.0 23.0 (+6.0)
DeepSeek-VL2-Small [51] 16.1B 28.3 44.1 (+15.8) 29.9 41.6 (+11.7) 23.6 26.0 (+2.4) 28.3 40.0 (+11.7)
Closed-source MLLMs (Full-Length Videos)
GPT-4o [39] - 57.6 64.8 (+7.2) 58.4 62.2 (+3.8) 43.3 46.5 (+3.2) 55.6 60.6 (+5.0)
o1 [38] - 72.4 73.0 (+0.6) 68.3 72.4 (+4.1) 52.0 65.4 (+13.4) 67.2 71.5 (+4.3)
Claude Sonnet 3.5 [5] - 7.9 17.4 (+9.5) 8.1 14.8 (+6.7) 7.9 17.3 (+9.4) 8.0 16.3 (+8.3)
Grok2-Vision [52] - 55.6 59.2 (+3.6) 55.8 55.2 (-0.6) 45.7 53.5 (+7.8) 54.1 56.5 (+2.4)
Nova-Lite [23] - 47.4 59.5 (+12.1) 45.3 55.9 (+10.6) 37.8 43.0 (+5.2) 44.9 55.3 (+10.4)
Nova-Pro [23] - 45.2 58.2 (+13.0) 48.4 58.1 (+9.7) 28.9 41.7 (+12.8) 44.1 55.5 (+11.4)
Gemini 1.5 Pro [49] - 65.1 69.4 (+4.3) 64.0 68.0 (+4.0) 54.3 55.1 (+0.8) 62.8 66.5 (+3.7)
Table 1: Performance comparison of different MLLMs across various settings with and without subtitles.
Video only Audio only Text only Video + Sub Video + Audio
Truncated Full Truncated Full Truncated Full Truncated Full Truncated Full
Level Easy 64.5 65.1 56.8 57.6 45.6 51.6 69.5 69.4 76.0 74.0
Medium 65.4 64.0 57.8 52.3 47.0 46.8 69.4 68.0 71.5 70.3
Hard 55.6 54.3 46.7 49.6 34.8 41.7 49.6 55.1 63.0 62.2
Agentic Tasks Social 55.6 58.3 47.5 49.0 41.4 54.2 62.6 59.4 61.6 67.7
Sentiment 69.1 70.4 53.1 51.9 34.6 42.0 66.7 64.2 76.5 74.1
Egocentric Agent 64.4 61.7 49.2 53.2 33.9 31.9 57.6 55.3 78.0 68.1
Information Querying 66.8 65.9 58.9 55.6 48.3 50.0 69.4 68.3 75.1 72.2
Sports 49.6 50.5 54.6 47.6 45.4 45.6 64.7 66.0 63.9 62.1
Gaming 63.1 61.2 53.4 49.4 41.7 43.5 58.3 64.7 63.1 62.4
Shopping 72.2 67.6 62.7 65.8 51.6 54.1 74.6 76.6 82.5 81.1
Audio Type Natural Sound 57.4 64.6 47.1 50.8 29.4 32.3 54.4 61.5 64.7 61.5
Speech 63.5 61.4 58.7 55.2 51.6 55.6 69.3 70.3 76.2 73.1
Music 69.2 72.0 53.8 56.0 7.7 24.0 69.2 56.0 76.9 76.0
Artificial Sound 60.0 74.0 36.0 40.0 20.0 26.0 56.0 58.0 56.0 68.0
Mixed Sounds 65.3 61.5 56.5 55.5 44.7 45.4 66.0 63.8 69.1 67.9
Topics Humanities and Society 70.7 71.3 63.9 65.9 48.9 51.9 77.4 79.1 82.7 80.6
Geography and Travel 68.8 68.8 56.2 56.2 62.5 68.8 50.0 68.8 75.0 75.0
Technology and Gaming 64.9 64.0 52.7 49.1 42.0 41.2 61.1 62.3 65.6 64.9
Science and Knowledge 57.1 63.0 39.3 51.9 39.3 37.0 53.6 63.0 67.9 70.4
Movies, TV and Animations 61.2 61.2 35.8 37.3 29.9 44.8 55.2 59.7 59.7 65.7
Arts and Performance 58.6 60.7 48.3 42.9 34.5 39.3 58.6 57.1 62.1 71.4
Pets and Animals 64.4 66.7 55.6 48.9 40.0 44.4 64.4 60.0 71.1 73.3
Business and Commerce 73.3 68.6 68.6 71.4 55.8 57.1 79.1 77.1 86.0 87.1
Life and Practical Skills 72.1 64.7 69.2 60.0 54.8 58.8 77.9 64.7 84.6 77.9
Sports and Adventure 53.2 56.9 52.4 48.6 43.7 45.9 65.1 69.7 63.5 63.3
Social Trends and Reactions 51.9 50.0 40.7 46.2 38.9 48.1 57.4 57.7 59.3 55.8
Vehicles and Transportation 55.6 55.6 55.6 44.4 22.2 11.1 55.6 44.4 66.7 55.6
Low Quality and Extended Content 41.7 41.7 54.2 54.2 29.2 37.5 37.5 50.0 66.7 62.5
Multimodal Abilities Visual Reasoning 63.2 65.7 54.8 57.4 43.4 46.6 64.5 64.2 72.8 71.1
Emotional Inference 67.4 68.3 55.8 51.2 37.2 34.1 62.8 58.5 72.1 73.2
Situational Reasoning 70.0 77.8 55.0 50.0 50.0 44.4 70.0 66.7 80.0 77.8
Causal Reasoning 65.7 60.3 47.8 43.1 37.3 48.3 67.2 65.5 70.1 60.3
Factual Recall 64.3 61.0 57.9 55.1 46.2 48.6 68.2 69.0 72.5 71.7
Spatial Reasoning 37.9 45.8 51.7 33.3 41.4 41.7 58.6 45.8 55.2 50.0
Relationship 62.5 62.5 43.8 62.5 50.0 75.0 62.5 75.0 62.5 75.0
Temporal Distance 66.7 81.8 58.3 54.5 58.3 72.7 58.3 81.8 83.3 81.8
Overall 63.5 62.8 55.6 53.9 44.5 47.9 66.3 66.5 71.9 70.5
Table 2: Performance comparison of Gemini 1.5 Pro across different modalities, difficulties, and categories.
We select a wide range of open-sourced and proprietary MLLMs on the MAVERIX benchmark. We report the quantitative results grouped by difficulties in Tab. 1 for their performances with and without subtitles and by different video lengths. The detailed quantitative GPT consistency scores can by found in Tab. D.2, and visualized in Fig. 6. We then showcase the fine-grained quantitative evaluations of Gemini 1.5 Pro in different multimodal input settings and grouped by the splits in Tab. 2. Refer to Section D for additional experimental results. Below, we provide our key observations.
Figure 6: Consistency of model responses on open-ended questions, measuring how reliably models provide coherent and relevant answers across different inputs.
@@ 3.3 Results
Are Audio and Text Equally Informative? Refer to Tab. 2 for detailed results. We evaluated Gemini 1.5 Pro, which was selected for its support for audio, text, and video modalities. Across all categories, Gemini 1.5 Pro consistently achieves higher scores with audio-only input compared to text-only input, suggesting that the model may extract more informative or discriminative cues from audio signals than from textual descriptions in this setting. Furthermore, when comparing multimodal configurations, video+audio yields the highest overall performance, surpassing video+subtitle.
Although video + audio generally provide the most significant benefit, certain categories, including sentiment understanding, situational reasoning, causal reasoning, and egocentric agent tasks, experience particularly large performance gains under this configuration. These tasks likely benefit from the richer contextual and temporal cues present in audio, such as tone, ambient sounds, and speaker intent, which are difficult to fully capture through visual or textual modalities alone.
Misaligned Category Trends Between MLLMs and Humans: The category-level analysis in Fig. 1 reveals notable differences in how MLLMs and humans perform across tasks. The quantitative comparisons are detailed in Tab. D.6. Humans tend to excel in dynamic contextual categories such as egocentric reasoning, sports analysis, and gaming strategy, which require interpreting spatial cues, tracking actions over time, and understanding complex interactions. However, their performance is less consistent in domains like shopping and social interaction. Shopping tasks may prioritize factual or procedural details that are less engaging or familiar to humans. For social interaction tasks, inherent subjectivity, such as interpreting cultural contexts, intentions, or unspoken tensions, can lead to diverse and sometimes conflicting human judgments, reducing overall precision.
MLLMs show a different trend. They perform comparatively well on shopping-related tasks, which often involve concrete retrieval of explicit audiovisual details like product descriptions, prices, or transactional cues. These structured object-centric queries align well with the strengths of large models trained on web-scale corpora rich in product and factual information. In contrast, MLLMs consistently struggle in categories that require nuanced temporal reasoning, social understanding, and dynamic context integration, such as predicting team strategies in games or recognizing subtle emotional cues in conversation.
These trends suggest that current MLLMs are more adept at tasks grounded in static, factual associations but remain limited in handling complex, situational reasoning that requires modeling interactions, intent, and evolving context. Closing this gap will likely require advances in temporal modeling, sentiment cue understanding, and the ability to generalize beyond explicit audiovisual patterns.
4 Related Work
MLLM Benchmarks: The development of MLLM benchmarks has progressed from static image-based question answering (QA) to dynamic video understanding. Early benchmarks for evaluating vision-language models (VLMs) focused primarily on images, emphasizing visual grounding and object recognition. Earlier datasets  [11, 3, 59, 41] provided human-annotated captions and QA pairs to assess these core capabilities. Building on this foundation, subsequent benchmarks [42, 34, 34, 44] introduced more domain-specific and knowledge-intensive tasks. A-OKVQA [43] challenges abilities to read and interpret texts within images. More recently, MMMU [54] and MMMU-Pro [55] have expanded these efforts with diverse tasks designed to probe a model’s ability to simultaneously ‘see” and read” from prompted images.
Despite their advances, image-based benchmarks lack the temporal and acoustic context necessary for richer multimodal understanding. This limitation motivated the shift toward video-based evaluation. Fine-grained temporal tasks assessing a model’s ability to perceive motion and temporal sequences have been included in video benchmarks [29, 16, 40, 25, 22, 45, 35, 30]. However, these datasets and recent extensions  [37, 12, 21, 36, 46] tend to feature short clips with constrained domains and limited semantic diversity. They primarily probe local temporal, visual semantics, and factual information, offering limited insight into how models handle the complexities of daily activities, situational awareness, or social reasoning. Most recent works [9, 29, 40, 25, 30, 10] take a step towards everyday contexts by incorporating human actions, but their scopes are restricted to human-centric activities and do not handle visual-audio cues. Video-MME [17] is a comprehensive video benchmark with short, medium, and long videos. However, it only employed the MCQ format, overlooking the quality and richness of free-form expressions generated by MLLMs from open-ended questions.
MAVERIX extends prior benchmarks by explicitly elevating audio as a primary modality, with its questions requiring complex cross-modal reasoning, such as resolving conflicts between dialogue and visual cues. Unlike existing datasets, it systematically evaluates a model’s ability to integrate complete multimodal signals through both descriptive open-ended answers as well as multiple-choice questions, comparisons shown in Tab. 3.
Video Understanding Models: Models based on contrastive pre-training over image-text pairs have demonstrated significant efficacy in learning transferable representations across various visual tasks. These methods have been adapted to specialized domains, including multi-disciplinary scientific understanding and multimodal reasoning. Building upon advancements in pre-training, subsequent work [50, 6, 48, 56, 57, 32] has sought to enhance the multimodal reasoning capabilities of complex spatio-temporal video data and improve their alignments with human preferences through post-training reinforcement learning from human feedback (RLHF). Other efforts [51, 31, 15, 47, 7] integrating a mixture of experts (MoE) architectures [33], which conditionally activates subsets of specialized “expert” networks during processing, have shown improved scalability and performance without a proportional increase in computational costs. However, whether these developments enable models to reason in ways comparable to or surpass humans remains an open question [1, 4, 8, 24], as the assessment of human-like multimodal reasoning, particularly in real-world visual-audio contexts, remains underexplored.
@@ 4 Related Work
Benchmarks #Videos Len. #Questions Subtitle Audio Open Multi-Split
MSRVTT-QA 2,990 15.2 72,821
MSVD-QA 504 9.2 50,505
TGIF-QA 9,575 3.6 165,165
ActivityNet-QA 5,800 14.9 58,000
TVQA 2,179 15.2 152,545
How2QA 1,517 11.5 71,812
STAR 914 11.9 59,871
NExT-QA 4,269 12.1 52,631
@@
Video-Bench 5,917 56.0 17,000
EgoSchema 6,365 18.0 36,054
AutoEval-Video 327 32.1 450
TempCompass 410 11.4 1,540
Video-MME 900 1071.9 2,700
HourVideo 500 2742 12,976
@@
MAVERIX 700 344.9 2,556
Table 3: Comparison of video QA benchmarks with various attributes.
5 Conclusion
Agentic tasks such as assisting collaborative work and navigating dynamic environments require strong audiovisual reasoning capabilities, and yet these abilities have not been fully assessed in the recent MLLMs. To address this, we introduce MAVERIX, a benchmark designed to evaluate models in complex real-world audiovisual scenarios. It contains 700 videos and 2,533 human carefully crafted questions providing not only a challenging benchmark but also a comprehensive evaluation suite. We conduct a thorough benchmarking of mainstream closed-source and open-source MLLMs on the proposed benchmark. Our evaluations reveal that cross-modal video-audio understanding is essential for solving real-world perception problems. Additionally, they show that even state-of-the-art MLLMs still exhibit a performance gap compared to humans, particularly in socially grounded or dynamic tasks. We hope MAVERIX will serve as a foundation for advancing research in multimodal reasoning and guiding future work towards more context-aware and socially intelligent systems.
Supplementary Material
MAVERIX Supplementary Material
  • Section A: MAVERIX Release v1.0
  • Section B: Rationale and Liminations
  • Section C: Qualitative Comparisons by Modalities
  • Section D: Additional Experiments
    • Section D.1: Standardized Prompt Design
    • Section D.2: Performance by Difficulties
    • Section D.4: Quantitative GPT Evaluations
    • Section D.5: Performance by Categories
    • Section D.6: Common Error Modes
  • Section E: Details on the Human Survey
  • Section F: Additional Data Distribution Analysis
Appendix A MAVERIX Release v1.0

### v2

@@ Introduction
All models receive inputs in the unified format [video frames, subtitles, question], with frames uniformly sampled at their maximum supported context window. For audio-incapable models, subtitles replace raw audio tracks, while the audio-supported models additionally process synchronized audio-video pairs. We employ a standardized prompt template across models, ensuring fairness by eliminating instructional biases. This setup isolates modality interdependence as the critical challenge: models must synthesize potentially asynchronous audiovisual cues, such as the startled expression of a character with an auditory context like an off-screen crash to match human-like understanding.
Table 4: Difficulty-wise MCQ accuracy (%) for audio-enabled models. AV cells show $\Delta$ vs A.
Model Easy Medium Hard
A V AV A V AV A V AV
@@
[…]
Model analysis with Maverix
This section discusses model behaviour along five dimensions: multimodal gains across modal deisgns, training‑recipe variation, agentic ability relative to humans, temporal‑horizon sensitivity, and perceived question difficulty, using the benchmark set.
Multimodal Gains Across Model Capacity and Architecture. Tab. 3 summarizes performance across five architectural families and multiple model sizes. Best unimodal accuracies span $\sim$26–55%, and most models show gains when additional modalities are provided. Many systems follow an audiovisual encoder++MLP++LLM design; among these, Ola, EgoGPT, and Qwen 2.5 Omni, which incorporate Whisper v3 for audio, generally do not regress and often improve relative to their strongest unimodal scores, whereas VITA 1.5 shows a regression with audio–visual input. The lightweight Gemini 2.0-FL and 2.5-FL variants also improve with multimodal inputs. For video-text models, adding a modality yields consistent gains; Qwen2-VL, Nova-Lite, Nova-Pro, and Claude 3.5 Sonnet improve by roughly $\sim$10%. Despite these gains, a sizable gap to human performance remains, suggesting that current models under-utilize cross-modal cues. This is evident when contrasting V++S and V++A: several models (e.g., Gemini 2.5-FL, EgoGPT, VITA 1.5, and Qwen 2.5 Omni) score lower with V++A than with V++S, indicating missed auditory details or limitations in audio-video fusion.
Table 5: Accuracy (%) on AV inputs across taxonomy and audio categories (single-column).
Taxonomy (AV)
Models Social Ego-centric Sports Gaming
Human 92.7 95.2 82.3 74.5
EgoGPT 34.3 45.8 42.9 40.8
Ola 52.5 57.6 52.9 46.6
VITA 1.5 21.2 6.8 16.8 12.6
Qwen 2.5 Omni 47.5 42.4 49.6 41.7
Gemini 2.0-FL 46.5 32.3 48.7 47.6
Gemini 2.5-FL 47.1 39.0 54.3 61.9
Audio Category (AV)
Models Natural Speech Music Artificial Mixed
Human 88.4 94.1 90.3 89.6 92.8
EgoGPT 41.2 44.2 46.2 48.0 47.3
Ola 48.5 52.0 38.5 58.0 56.5
VITA 1.5 29.4 17.0 7.7 26.0 17.6
Qwen 2.5 Omni 52.9 47.8 53.8 48.0 51.5
Gemini 2.0-FL 47.1 51.1 38.5 54.0 49.6
Gemini 2.5-FL 51.5 56.7 50.0 52.0 53.1
Training-Recipe Variants: SFT, RL, and Data Composition. We assess open-source recipes built on identical backbones to isolate curriculum effects. Omni-modal models such as Ola and Qwen 2.5 Omni use an image-text warmup followed by separate alignment for the audio modality, and they typically improve on audiovisual evaluations; reinforcement learning appears to further increase the multimodal gains. VITA 1.5, on the other hand, emphasizes alignment to the video modality during training, which may bias attention toward the visual stream and results in regression when subtitles and audio are added.
Video-text models follow a more streamlined path: initial pretraining on image-caption pairs to align images with text, then instruction tuning or long chain-of-thought data for fine-tuning. Aside from minor architectural differences, they vary primarily in data curation and sources. Qwen 2.5 VL uniquely includes chain-of-thought data during fine-tuning to encourage explicit reasoning and stronger multimodal synthesis. While its overall performance is strong, the relative gain from multimodal inputs appears similar to its counterpart without chain-of-thought fine-tuning, which may point to reward hacking during SFT and warrants further study.
Table 6: Open-ended response correctness scores (out of 5) as judged by GPT, reported per modality. For models without native audio support, scores for Subtitle (S) and Subtitle+Video (SV) are shown instead of Audio (A) and Audio+Video (AV). Response token length statistics are analyzed separately.
Model
Audio
Support
Judged Score Token Length
A/S V AV/SV Avg Max Std
[…]
Qwen-2-Omni 1.0 1.2 1.2 $0\downarrow$ 47.0 178 26.5
[…]
Agentic Ability in Comparison with Humans. Tab. 5 indicates that humans perform best on social and egocentric questions, with slightly lower accuracy on gaming and sports that demand domain knowledge. Model behavior is less uniform. Gemini, Qwen 2.5 Omni, and VITA 1.5 tend to be weaker on egocentric videos and comparatively stronger on domain-specific categories, while the remaining models show different per-category strengths.
Across audio categories, human accuracy is largely stable. Models trained on broader multimodal corpora such as Qwen 2.5 Omni and the Gemini family exhibit smaller fluctuations across audio types, yet many systems under-perform when music dominates. EgoGPT shows relatively strong auditory understanding, likely reflecting its use of a pretrained Whisper encoder. In contrast, Ola and VITA 1.5 that trained on smaller datasets, display larger variance across categories, with notable drops on the music category.
Taken together, these patterns suggest that coverage of first-person content and diverse audio during training may be as important as scale for robust agentic ability across domains.
Temporal Horizons: Long- versus Short-Clip Performance. As shown in Fig. 4, we evaluate models on short, pre-localized audiovisual clips in MAVERIX and on their full-length counterparts in MAVERIX-Long. Across models, localized clips yield higher accuracy. Among the agentic categories, the questions that depend on immediate, synchronous audiovisual cues, such as those from factual recall and near-term causal inference, show the smallest degradation. When the relevant segment is pre-localized, models can more reliably extract the necessary information.
By contrast, social relationship, emotion, and situational understanding often rely on fine-grained and sometimes asynchronous cues distributed over time. Performance drops more on long videos, reflecting challenges in localizing these signals and integrating them over extended context. Overall, a gap to human performance remains, especially for longer videos and for recognizing subtle contextual cues. These trends suggest that current MLLMs are stronger at retrieving salient, object or event level signals than at integrating evolving context and social nuance over time.
Figure 4: Impact of Video Length on Agentic Category Performances. (a) Accuracy change (%) from short to full-length videos across models and questions. (b) Accuracy gap (%) to human performance for short and full inputs.
Model and Human Perception of Difficulty. We analyze performance by difficulty and observe that multimodal inputs often help most on easy items, with smaller gains on hard ones, though trends vary by model. For the audio-enabled models where split statistics are available, Gemini 2.0-FL improves by ++12.4% on easy items and ++3.3% on hard items, and Gemini 2.5-FL improves by ++11.5% on easy items and ++5.3% on hard items. Qwen 2.5 Omni shows a different pattern with substantial benefit on hard items as well. These mixed results suggest that current systems leverage straightforward cross-modal cues more reliably than they integrate sparse or subtle signals in harder cases.
Although MAVERIX is designed to elicit cross-modal reasoning, some models still achieve moderate scores with a single modality, likely because many real videos contain aligned audio and visual streams that allow plausible inferences from partial evidence. GPT-4o-mini is one such example of respectable unimodal performance. Humans also benefit from aligned cues and strong priors, yet the jump from 81.4% with video-only to 92.8% with audiovisual highlights the value of genuine cross-modal understanding and sets a clear target for future modeling. We hope continued progress in cross-modal alignment will narrow this gap and eventually surpass the current human baseline.
Related Work
MLLM Benchmarks. Early vision-language benchmarks centered on images for grounding and recognition, using captions and QA pairs (Chen et al. 2015; Agrawal et al. 2019; Zhu et al. 2016; Plummer et al. 2015), followed by domain-specific and knowledge-intensive settings (Saikh et al. 2022; Lu et al. 2023; Singh et al. 2019). A-OKVQA targets external-knowledge reasoning beyond visible content (Schwenk et al. 2022). More recent efforts, including MMMU and MMMU-Pro, broaden question diversity and reading-from-image skills (Yue et al. 2024a, b).
Image-only evaluation lacks temporal and acoustic context, motivating video benchmarks that probe motion, sequence, and temporal localization (Li et al. 2024b; Patraucean et al. 2023; Kesen et al. 2023; Huang et al. 2020; Song et al. 2024a; Maaz et al. 2024; Fang et al. 2025; Li et al. 2024c; Ning et al. 2023; Chen et al. 2024a; He et al. 2024; Mangalam, Akshulakov, and Malik 2023; Song et al. 2024b). However, most emphasize short clips and constrained domains, rely on MCQ-only protocols, and provide limited coverage of everyday social or situational reasoning. Video-MME and AV-Odyssey scale video duration but remain MCQ-only, omitting open-ended assessment (Fu et al. 2024; Gong et al. 2024). MAVERIX elevates audio as one of the primary signals and stresses cross-modal integration as it evaluates both 8-way MCQs and open-ended responses to measure multimodal synthesis under realistic audiovisual conditions.
Video Understanding Models. Contrastive pretraining on image-text data yields transferable representations and has been adapted to spatio-temporal reasoning; post-training with instruction tuning and RLHF further aligns models with human preferences (Sun et al. 2023; Zhai et al. 2024; Lin et al. 2023; Wang et al. 2024; Bai et al. 2025; Zhang 2024). MoE-style routing improves scalability without linear cost growth (Wu et al. 2024; Lin et al. 2024; DeepSeek-AI and Aixin Liu 2025; Sun, Chen, and Yiqing Huang 2024; Cai et al. 2024; Liu et al. 2025a). Yet whether these advances enable human-comparable multimodal reasoning in real-world audiovisual settings remains open (tes 2024; Amirizaniani et al. 2024; Campbell et al. 2024; Kazemi et al. 2024). Our evaluations on MAVERIX show a substantial gap to human accuracy on MCQs, often on the order of several tens of percentage points, and highlight persistent challenges in integrating temporal, social, and auditory cues in models with different architecture and training recipes.
Conclusion
Agentic scenarios such as assisting collaborative work and navigating dynamic environments require strong audiovisual reasoning, yet these abilities remain under-assessed in recent MLLMs. We introduce MAVERIX, a benchmark for complex, real-world audiovisual understanding, comprising 700 videos and 2,556 carefully crafted, human-authored questions. The suite evaluates models with both 8-way multiple-choice and open-ended responses.
Our results indicate that multimodal inputs generally improve accuracy, but sizable gaps to human performance persist, especially for socially grounded or dynamic scenarios. Models benefit most when relevant segments are pre-localized and tend to struggle on longer videos that demand integrating subtle, asynchronous cues over time; robust audio integration also remains uneven across systems. We hope MAVERIX will guide progress toward stronger cross-modal alignment, better temporal reasoning, and more context-aware, socially intelligent models.
Appendix A A. MAVERIX Release v1.0